OpenStudy (anonymous):

Could someone please help? I am sooo confuzzled :( -4√(x+9)=20

OpenStudy (campbell_st):

start by dividing both sides of the equation by -4 what do you get..?

OpenStudy (anonymous):

sqrt(x+9)=5

OpenStudy (anonymous):

Right?

OpenStudy (campbell_st):

great now square both sides of the equation and you made a slight error... which won't affect the answer it should be \[\sqrt{x + 9} = -5\] and remember when you square a negative you get a positive number..

OpenStudy (anonymous):

Why is the 5 negative though?

OpenStudy (anonymous):

Oh! Because the 4 is negative! I was looking at it wrong.

OpenStudy (campbell_st):

its -5 because you divided a positive number by an neative number 20/(-4) = -5

OpenStudy (anonymous):

Thanks! I think I get it now. Is there more to work out next? Or am I just left with sqrt(x+9)=-5?

OpenStudy (campbell_st):

ok... what did you get when you squared both sides of the equation..?

OpenStudy (campbell_st):

this is a 3 step equation... step 1: divide by -4 step 2: square both sides

OpenStudy (anonymous):

sqrt(x+81)=25?

OpenStudy (campbell_st):

slight error... \[(\sqrt{x + 9})^2 = x + 9\] the reason is to do with the index laws a square root is a power of 1/2 so its \[((x + 9)^{\frac{1}{2}})^2\] and the power of a power rule says \[(x^a)^b = x ^{a \times b}\] so you need to multiply the powers 1/2 x 2 = 1 so its (x + 9)^1 which means you have x + 9 = 25 just solve for x

OpenStudy (mertsj):

There is one slight problem here: At the very beginning when you divided both sides by -4 you got: \[\sqrt{x+9}=-5\]

OpenStudy (mertsj):

Now that equation says that the POSITIVE square root is equal to a NEGATIVE number. So you know right off that there is NO SOLUTION to this problem.

OpenStudy (mertsj):

And that 16 is an extraneous root

OpenStudy (anonymous):

Yes?

OpenStudy (mertsj):

So the answer is that there is no answer.

OpenStudy (anonymous):

Oh. So do I stop after dividing or do I stop after squaring both sides?

OpenStudy (anonymous):

I would write it like this... -4√(x+9)=20 -4√((x)+9)/4=20/4 √(x+9)2=52 √(x+81)=25 No solution

OpenStudy (mertsj):

You could stop after dividing is you noticed the problem or you could proceed to the end and find that you get an answer of 16. Then check it and when you find it won't check discard it and declare no solution.

OpenStudy (mertsj):

Did your teacher teach you that you must always check a solution that comes from a radical equation because sometimes the squaring process introduces an extraneous root.

OpenStudy (anonymous):

No?

OpenStudy (mertsj):

So now you have been taught. Always check the answers to radical equations. Sometimes they will not check.

OpenStudy (anonymous):

Okay